Let A be an m x r matrix, B be an m x s matrix, C be an r x n matrix, and D be an s x n matrix.
Prove that
(A B) where A and B are side by side in the 1 x 2 bracket, multiply (C D) where C is on top of D in the 2x1 bracket = AC + BD
1
Expert's answer
2020-02-17T08:20:26-0500

We are given that there is one matrix of order "1\times2" "\ \ \begin{pmatrix}
 a & b
 
\end{pmatrix}" and other matrix of order 2"\times"1 "\begin{pmatrix}
 c \\
 d
\end{pmatrix}"

We have to prove that their product sum is equal to ("ac+bd" )

To multiply matrices, the number of columns of one of them must be equal to the number of rows of another one , here this condition is fulfilled and their product will be a 1x 1 matrix

let's multiply

"\begin{pmatrix} a & b\end{pmatrix}\cdot \begin{pmatrix} c\\ d \end{pmatrix}""=a*c+b*d"

We obtained this result.